# Modelling Mean-Field Games with Neural Ordinary Differential Equations

### v2

@@ 1. $j_{1}$
  2. $j_{2}$
  3. $j_{3}$
$j_{1}(s,\tilde{\tau})$ $=\left[\tilde{\tau}-s,0\right]_{+}$
$j_{2}(\tilde{s},\tilde{\tau})$ $=[\tilde{\tau}-\tilde{s},0]_{+}$
$j_{3}(\tilde{s},\tilde{\tau})$ $=[\tilde{s}-\tilde{\tau},0]_{+}$
Figure 2: The distribution of meeting arrival times for the standard (a) and the neural ODE (b) version of the meeting arrival times game. In (a), the randomly initialized distribution of arrival times quickly converges to a narrow distribution centred at $\tilde{s}$. The value of $\tilde{s}$ is subject to the Brownian noise of the SDE. In (b), the distribution of meeting arrival times of agents whose control is steered towards arriving at 15:00 in turn 15.
The behaviour of all other agents is summarized in the actual starting time $\tilde{s}$, which is the mean-field of this game by summarizing all the agents’ actions in a cumulative distribution of arrival times. To acknowledge the coupling between $\tilde{s}$ and the individual agents, consider that each agent bases their intended arrival $\tau^{i}$ time on $\tilde{s}$ and $\tilde{s}$ is influenced by the cumulative distribution of all actual arrival times $\tilde{\tau}^{i}$. In addition, Guéant et al. prove that $\tilde{s}$ exists, being a unique fixed point representing the game’s equilibrium.
We model the “standard" version of the game with a SDE that differentiates over the turns $t$ (see Equation 1). Even though the game lacks a terminal condition, we treat it as a finite game by setting $t\in[1,15]$. Figure 2a shows an example outcome of the game, where the distribution of arrival times quickly narrows down. This observation is in accord with the proof provided by Guéant et al. The exact value of $\tilde{s}$ diffuses between turns due to the Brownian motion in the SDE. While the original game description does not include Brownian motion, it is included in this example to demonstrate the modelling capacity of the neural ODE.
4.1.1 Learning the intended arrival times
[…]
4.2 The El Farol Bar Problem
Even though the meeting time arrival game provides a proof of concept, the neural ODE does not yield additional insights into the game dynamics. Therefore, we consider an implementation of the El Farol Bar problem, a more complex example introduced by Arthur [40]. In this problem, a fixed population of agents can go to a bar. Each agent $i$ goes to the bar with a probability $p_{i}$, and all agents need to decide simultaneously whether they’re going. The agents’ utility depends on the bar’s crowdedness. If the bar is too crowded, all the agents that went there regret their decision and have less fun than the ones that stayed at home.
The cost function $J$ associated with the bar problem is comparable to that of the meeting arrival times game. However, the new cost function is selective; not all components apply to all agents. This selectivity is presumed to lead to an interplay between the MFG dynamics and the neural network. The cost is parameterized by the agent’s probability of going to the bar ($p$), the fraction of agents that went to the bar ($a$), and the crowding threshold ($c$). $J$ consists of the sum of $j_{[1-3]}$ described below.
  1. $j_{1}$
    Missed a good evening The cost associated with staying at home while the bar was not too crowded. This cost only applies to agents that did not go to the bar.
  2. $j_{2}$
    Crowding penalty The cost associated with going to a full bar. This cost only applies to the agents that went to the bar.
  3. $j_{3}$
$j_{1}(p,a,c)$ =$=\begin{cases}0&\text{if }a\geq c\\ [c-p]_{+}&\text{if }a<c\\ \end{cases}$
$j_{2}(p,a,c)$ =$=\begin{cases}[p-c]_{+}&\text{if }a\geq c\\ 0&\text{if }a<c\end{cases}$
$j_{3}(p,a)$ $=\left(p-a\right)^{2}$
In contrast with the meeting arrival times game, the value function of the El Farol Bar problem does not include a Brownian motion. Instead, it is modelled as an ODE differentiating over $t\in[1,15]$, where each turn presents a new opportunity to go to the bar. As the distribution of agents is summarized in the fraction of agents that went to the bar, $a$ can be considered the mean-field of the game. The agents reach a Nash equilibrium when the mean probability of going to the bar converges to the crowding threshold [40]. This behaviour is highlighted in Figure 3a. The figure shows the evolution of the distribution of agents, where $p$ goes from a uniform distribution on the interval $\left[0.0,0.1\right\rangle$ at turn 1, to a narrow distribution centred around $c=0.9$ at turn 15.
4.2.1 Observed attendance rates and emerging strategies
Whereas the dynamics observed in Figure 3a purely emerge from the game’s setup, we test whether a combination of the game rules with observed attendance rates leads to more informed strategies. To generate observations, we consider $N\times M$ agents. To create a single data point $\bm{x}\in\mathbb{R}^{10}$, the agents’ attendance rates are monitored for 10 subsequent turns. The attendance rate of one evening is generated by summing over $j\in[1,2,\dots,N\times M]$ samples drawn from $Bern\left(p_{j}\right)$. Just like the arrival times data, the vector of probabilities is created by drawing $M$ samples from $N$ different Gaussians $\mathcal{N}\left(\mu_{i},\sigma_{i}\right)$ for $i\in[1,2,\dots,N]$. The hyperparameters are the same for both games. However, the agents’ intentions $p_{j}$ are limited to the domain $\left[0,1\right]$ to represent a probability.
To discover how the agents’ probabilities of going to the bar change after training on 10 observations in $\mathbb{R}^{10}$, we deliberately diverge from the MFG setup by distributing the observed attendance rates around $\mu_{0}=0.2$ (Figure 3b, red). Note that $\mu_{0}$ is significantly below the crowding threshold $c=0.9$ (Figure 3b, blue). Even though the agents can only decide on a single output value, the neural ODE allows the agents to build more complex strategies than the standard MFG. Namely, the MFG-based agents in Figure 3a converge to an almost homogeneous strategy $p=c$. In contrast, the results from Figure 3b are more widely distributed. Furthermore, the distribution “oscillates" between turns. This oscillation results from the interplay between $j_{1}$, drawing the agents towards the crowding threshold, and the combination of the neural network contribution and $j_{3}$, pulling the agents back to $\mu_{0}$. Regardless of the oscillations, there are a few agents with $p=1.0$, always going to the bar, independent of the other strategies. The agents from the last category leverage the fact that the majority of agents have a low probability of going to the bar. Because of this, they can always go to a bar that is never too crowded.
Figure 3: The distribution of probabilities of going to the bar for the standard (a) and neural ODE (b) version of the El Farol Bar problem. In the standard version, the distribution of probabilities converges towards the crowding threshold of $c=0.9$ (blue). The neural ODE (b) is trained to align with the observed attendance rates, where the mean is indicated in red. The density histograms show the normalized distributions at turn 15.
4.3 Liar’s Dice
The last game discussed is Liar’s dice, a dice game that starts with $M$ players rolling $D$ dice with $N$ faces each. Next, the players take turns, estimating the number of occurrences of a specific face in all dice rolls. The first player makes an opening bid, e.g. stating that there are 2 dice with face 4. Every subsequent player may bid a higher quantity of dice with the same face; a higher quantity of dice with a higher face; or the same quantity of dice with a higher face. Alternatively, a player may challenge the previous bid, ending the round. If the combined number of dice is smaller than the bidding quantity, the challenge is successful and the challenging player wins the round. The challenge is unsuccessful if the combined number of dice is larger than or equal to the bidding quantity, in which case the player who made the last bid wins.
In this partially observable game, the players aim to make deceitful, but probable bids. Therefore, the players’ utility is based on (1) their ability to bluff and (2) the correctness of their estimate of the dice odds. As it is very hard to capture these ideas in an appropriate cost function, we model $J$ as a starting point. We deliberately leave room for the neural network to learn the missing details leading to the optimal control. In short, $J$ is the sum of $j_{1}(q_{t},q_{t-1},m_{q})=[m_{q}-(q_{t}-q_{t-1})]_{+}$, $j_{2}(q_{t},a)=[a-q_{t}]_{+}$ and $j_{3}(f_{t},m_{f})=[m_{f}-f_{t}]_{+}$, explained below.
  1. $j_{1}$
@@
    Safety penalty A penalty for being on the safe side when increasing the bid’s quantity, making it easier for the next player to make a plausible bid. A metric of the quantities of the current ($q_{t}$) and previous ($q_{t-1}$) bid, and the maximal possible quantity of the face of interest ($m_{q}$).
@@ 3. $j_{3}$
    Niceness penalty A penalty for not increasing the face of the bids, leaving more room for the other players to make a plausible bid. A metric of the largest face of the die $m_{f}$ and the face associated with the current bid $f_{t}$.
@@
In each round, we draw $D$ dice throws from a categorical distribution parameterized by the true dice odds $\theta$. Next, the player’s bid is based on the MFG drift and the neural network output. The interested reader will find the pseudocode for the drift in Appendix A. The neural network output differs per turn $t$ and depends on $\hat{\theta}_{t}^{i}$ of the current player, the current player’s dice outcomes and the previous player’s bid. The neural network parameters are only updated when the round has ended. The loss fuelling this update consists of the mean squared error between the $\hat{\theta}_{t}^{i}$ and the distribution estimated from all dice outcomes revealed at the end of the round. In addition, the loss includes a term based on the correctness of the challenge. This term adds the batch-wise ratio of correct challenges to the loss, penalizing the network if a game ends with a correct challenge and optimizing $\lambda$.
4.3.1 Informed Bluffing
Figure 4a shows the influence of the initial bluff strategy on the game length. The results from Figure 4a suggest that the strategy from the neural ODE players (orange) tends to yield shorter games for $\lambda_{0}<3$ compared to the strategy based on the MFG dynamics (blue). This effect is caused by the larger drift of $\lambda$ in the neural ODE, where the neural ODE-based players are more likely to learn larger values for $\lambda_{t}$ for $t>0$ when $\lambda_{0}$ is small (Figure 4b). Despite the more risky bluffing strategy of the neural ODE strategy, there is no strong evidence of an increase in the number of successful challenges. The relation between the bluffing strategies and the ratio of successful challenges is further explored in Appendix B.
4.3.2 Estimating the Dice Odds
We test the efficiency of the model by testing its resilience against an incorrect belief $\hat{\theta}_{0}$. To do so, we simulate a new series of games, where $N=2$ agents unknowingly play with $D=2$ (the same) unfair dice, i.e. $\hat{\theta}_{0}\neq\theta$. The agents can only update their beliefs when a game ends and all four dice are revealed to verify the challenge. Given the loss function, it is expected that the more dice the agents see, the likelier they are to update their beliefs. This behaviour is summarized in Figure 5, which relates the Kullback-Leibler divergence between the true and estimated dice odds to the number of dice seen during training. [^2] Because the KL divergences are so small, we ignore the loss component regarding the bluff parameter for clarity. Even though the model is not solving a difficult task, it is still very effective and converges after seeing around 15 dice.
5 Discussion
Neural ODEs form a powerful tool for influencing the Nash equilibria of a game. Our experimental results highlight the model’s strengths: the results found with meeting arrival times game underline its ability to learn abstractions from data, even under noisy circumstances. In addition, the analyses of the El Farol bar problem exemplify strategies that emerge from the interaction between the data-driven dynamics and the MFG mechanics. Lastly, the neural ODE can learn more elaborate strategies without training on external observations. When training the neural ODE in the context of Liar’s dice, we obtain new bluffing strategies when applying the neural network to observations made in previous game states.
Limitations   A limitation of our methods is that the neural network introduces a “black box" into the game dynamics. In particular, the inclusion of the neural network reduces the game’s general interpretability by making the agents’ decisions more opaque. Furthermore, the formulation of the model is limited by the symmetry assumption of the MFG. Because of the large number of players involved, the neural network has to be shared between agents. While this does not imply that the agents have identical beliefs, they are still linked through the neural network parameters.
6 Conclusion
Mean-field game theory is used to efficiently model games with a very large to infinite population of agents. The agents’ Nash equilibria are commonly found via reinforcement learning or by solving the system of PDEs associated with the MFG. Both methods have their advantages, the first being model-free and the second numerically exact. To get the best of both worlds, we extend the PDE formulation of the MFG with a neural network. The resulting neural ODE exerts a data-driven influence on the game’s Nash equilibria, deriving strategies that are in line with observed game outcomes. In addition, it can be trained to recover structure from the game’s past dynamics, learning more complex strategies without using extra data. Leveraging those properties, the model can fill in dynamics that are unaccounted for in the original MFG specification. Thus, the addition of the neural network not only makes it easier to model MFGs, but also decreases the modelling bias compared to fully specified systems of PDEs. Besides their ability to incorporate observations, neural ODEs are a numerically exact and lightweight alternative to (deep) RL-based methods for solving MFGs. To match the RL-based methods in their flexibility, the neural ODE can be made completely model-free by fully omitting the predefined partial derivatives. Lastly, their dependency on automatic differentiation makes the neural ODEs more objective than approaches that solve PDEs using finite differences [28, 29], which may prevent the loss of existence or uniqueness of solutions associated with solving MFGs [15].
Even though we only consider simple, synthetic games, we hope that our work provides sufficient proof of concept to inspire future research on data-driven modelling of MFGs. For example, neural ODEs could be used to formulate poker, as has been done with deep counterfactual regret minimization by Brown [42]. Alternatively, the neural ODEs can be elaborated on by using a fixed number of neural networks to model subpopulations of agents. Such models could find applications in settings used to analyse collaborative or adversarial strategies. In addition to solving more complex games, we expect the model to find applications in systems associated with non-atomic, anonymous players and observations, which are commonly found in markets and economics [43].
Software and Data
The source code associated with the methods and experiments is available on GitHub: https://github.com/KachmanLab/NeuralMFG. All software was prepared by the Radboud University authors.
Appendix A Liar’s logic
The players of the Liar’s dice game need to decide at each turn whether they increment or challenge the previous bid. To do so, they follow the logic presented in Algorithm 1.
@@ Appendix A Liar’s logic
Appendix B Additional Liar’s dice analyses
B.1 The ratio of successful challenges
@@
As mentioned in Section 4.3.1, a more deceitful playing style of the neural ODE-based players does not seem to increase the ratio of successful challenges. To support this claim, we reproduce Figure 4b and compare it to the ratio of successful challenges presented in Figure 6b. When considering the ratio of successful challenges for $\lambda_{0}\leq 6$, the neural ODE-based agents (orange) have a smaller ratio than the purely MFG-based agents (blue). The only exception is found at $\lambda_{0}=0$, where the ratio of successful challenges for the MFG-based agents is very low due to the total absence of bluffing (Figure 6a). However, when $\lambda_{0}>6$, the success ratio of the neural ODE agents is larger than that of the MFG-based agents, even though the strategies of both groups are almost equal (Figure 6a).
@@
Figure 6: A comparison between the bluffing strategies of the agents that are purely based on the MFG mechanics (blue) and the agents that are based on a neural ODE (orange). All agents have a correct belief about the dice odds. (a) presents the learned versus initial bluff strategy and the ratio of successful challenges as a function of the initial bluff strategy is shown in (b). The results are aggregated over 100 games. The error bars indicate a standard deviation from this average.
B.2 Playing with unfair dice
@@
When analyzing the bluff strategies in the main text and Appendix B.1, we assume that the players’ beliefs about the dice are correct. In this section, we do not make this assumption and consider the interaction between the bluff strategy and the players’ beliefs about the dice odds. To do so, we again initialize a game with $N=30$ players having $D=15$ dice each. However, the players have the uniform initial beliefs that the dice are fair: $\hat{\theta}_{t}^{t}=(\nicefrac{{1}}{{6}},\nicefrac{{1}}{{6}},\nicefrac{{1}}{{6}},\nicefrac{{1}}{{6}},\nicefrac{{1}}{{6}},\nicefrac{{1}}{{6}})$. In contrast, the actual dice odds are $\theta=(\nicefrac{{1}}{{10}},\nicefrac{{2}}{{10}},\nicefrac{{0}}{{10}},\nicefrac{{4}}{{10}},\nicefrac{{1}}{{10}},\nicefrac{{2}}{{10}})$.

### v3

@@ 1. $j_{1}$
    Safety penalty A penalty for being on the safe side when increasing the bid’s quantity, making it easier for the next player to make a plausible bid. A metric of the quantities of the current ($q_{t}$) and previous ($q_{t-1}$) bid, and the maximal possible quantity of the face of interest ($m^{N}_{q}$).
    $j_{1}(q_{t},q_{t-1},m_{q}(t))=[m_{q}(t)-(q_{t}-q_{t-1})]_{+}$
@@ 3. $j_{3}$
    Niceness penalty A penalty for not increasing the face of the bids, leaving more room for the other players to make a plausible bid. A metric of the largest face of the die $m^{N}_{f}$ and the face associated with the current bid $f_{t}$.
    $j_{3}(f_{t},m_{f}(t)=[m_{f}(t)-f_{t}]_{+}$
@@
In each round, we draw $D$ dice throws from a categorical distribution parameterized by the true dice odds $\theta$. Next, the player’s bid is based on the MFG drift and the neural network output. The pseudocode describing the players’ drift is provided in Algorithm 1. The neural network output differs per turn $t$ and depends on $\hat{\theta}_{t}^{i}$ of the current player, the current player’s dice outcomes and the previous player’s bid. The neural network parameters are only updated when the round has ended. The loss fuelling this update consists of the mean squared error between the $\hat{\theta}_{t}^{i}$ and the distribution estimated from all dice outcomes revealed at the end of the round. In addition, the loss includes a term based on the correctness of the challenge. This term adds the batch-wise ratio of correct challenges to the loss, penalizing the network if a game ends with a correct challenge and optimizing $\lambda$.
@@
C.1 Informed Bluffing
Figure 9a shows the influence of the initial bluff strategy on the game length. The results from Figure 9a suggest that the strategy from the neural ODE players (orange) tends to yield shorter games for $\lambda_{0}<3$ compared to the strategy based on the MFG dynamics (blue). This effect is caused by the larger drift of $\lambda$ in the neural ODE, where the neural ODE-based players are more likely to learn larger values for $\lambda_{t}$ for $t>0$ when $\lambda_{0}$ is small (Figure 9b). Despite the more risky bluffing strategy of the neural ODE strategy, there is no strong evidence of an increase in the number of successful challenges. The relation between the bluffing strategies and the ratio of successful challenges is further explored in Section C.4.
C.2 Estimating the Dice Odds
We test the efficiency of the model by testing its resilience against an incorrect belief $\hat{\theta}_{0}$. To do so, we simulate a new series of games, where $N=2$ agents unknowingly play with $D=2$ (the same) unfair dice, i.e. $\hat{\theta}_{0}\neq\theta$. The agents can only update their beliefs when a game ends and all four dice are revealed to verify the challenge. Given the loss function, it is expected that the more dice the agents see, the more likely they are to update their beliefs. This behaviour is summarized in Figure 10, which relates the Kullback-Leibler divergence between the true and estimated dice odds to the number of dice seen during training. [^4] Because the KL divergences are so small, we ignore the loss component regarding the bluff parameter for clarity. Even though the model is not solving a difficult task, it is still very effective and converges after seeing around 15 dice.
C.3 The ratio of successful challenges
@@
As mentioned in Section C.1, a more deceitful playing style of the neural ODE-based players does not seem to increase the ratio of successful challenges. To support this claim, we reproduce Figure 9b and compare it to the ratio of successful challenges presented in Figure 11b. When considering the ratio of successful challenges for $\lambda_{0}\leq 6$, the neural ODE-based agents (orange) have a smaller ratio than the purely MFG-based agents (blue). The only exception is found at $\lambda_{0}=0$, where the ratio of successful challenges for the MFG-based agents is very low due to the total absence of bluffing (Figure 11a). However, when $\lambda_{0}>6$, the success ratio of the neural ODE agents is larger than that of the MFG-based agents, even though the strategies of both groups are almost equal (Figure 11a).
@@
Figure 11: A comparison between the bluffing strategies of the agents that are purely based on the MFG mechanics (blue) and the agents that are based on a neural ODE (orange). All agents have a correct belief about the dice odds. (a) presents the learned versus initial bluff strategy, and the ratio of successful challenges as a function of the initial bluff strategy is shown in (b). The results are aggregated over 100 games. The error bars indicate a standard deviation from this average.
C.4 Playing with unfair dice
@@
When analyzing the bluff strategies in the main text and Appendix C.3, we assume that the players’ beliefs about the dice are correct. In this section, we do not make this assumption and consider the interaction between the bluff strategy and the players’ beliefs about the dice odds. To do so, we again initialize a game with $N=30$ players having $D=15$ dice each. However, the players have the uniform initial beliefs that the dice are fair: $\hat{\theta}_{t}^{t}=(\nicefrac{{1}}{{6}},\nicefrac{{1}}{{6}},\nicefrac{{1}}{{6}},\nicefrac{{1}}{{6}},\nicefrac{{1}}{{6}},\nicefrac{{1}}{{6}})$. In contrast, the actual dice odds are $\theta=(\nicefrac{{1}}{{10}},\nicefrac{{2}}{{10}},\nicefrac{{0}}{{10}},\nicefrac{{4}}{{10}},\nicefrac{{1}}{{10}},\nicefrac{{2}}{{10}})$.